Solve the following System of Linear Equations using the Elimination Method. 2x-2y=6
4x+2y=18
Answer and show all your work.

Found 2 solutions by Solver92311, MathTherapy:
Answer by Solver92311(821)



Multiply the first equation by -2. Then add the like terms which will eliminate the 'x' term leaving an equation in 'y' only. Solve the resulting equation in 'y'. Substitute the resulting value of 'y' into either of the original equations to create an equation in 'x'. Solve for 'x'. The answer should be an ordered pair (x,y).


John

My calculator said it, I believe it, that settles it

From
I > Ø

Answer by MathTherapy(10808)


Solve the following System of Linear Equations using the Elimination Method. 2x-2y=6
4x+2y=18
Answer and show all your work.
2x - 2y = 6 ---- eq (i)
4x + 2y = 18 -------- eq (ii)
ADDING the 2 equations ELIMINATES y. Solve for x to get its value.
Substitute the value of x in any of the 2 equations to get the value of y.
"Answer and show your work." Are you serious? Whose job do you think that is?
You've been given instructions, now do the REST!!